OpenStudy (anonymous):

Write the equation of the line that passes through (3, 6) and (0, –3) in slope-intercept form. y = 3x – 6 y = 3x – 3 y = 3x + 3 y = 3x + 6

OpenStudy (jdoe0001):

what would you say is the slope of THAT line?

OpenStudy (anonymous):

y=mx+c is the formula to use ..... m=y2-y1/x2-x1

OpenStudy (anonymous):

for c /.... compare ....\[m=\frac{ y-c }{ x-o }-----with ------m=\frac{ y2-y1 }{x2-x1}\]

OpenStudy (anonymous):

so the answer is c?

OpenStudy (anonymous):

hmmm ....

OpenStudy (anonymous):

lets find the slope... slope(m) = (y2 - y1) / (x2 - x1) set 1 (3,6) set 2 (0,-3) Now lets sub it in... m = (-3 - 6) / (0 - 3) m = - 9/-3 m = 3 (remember when dividing 2 negatives it equals a positive) so now we know the slope is 3 Now we will use this formula, y = mx + b.....using 3 as slope m and either of the points.....I will use (3,6) now lets sub... 6 = 3(3) + b --(we are solving for b, the y intercept) 6 = 9 + b 6 - 9 = b - 3 = b so your equation is : y = 3x - 3